# Better wit than wealth: Dynamic Parametric Retrieval Augmented Generation for Test-time Knowledge Enhancement

### v1

@@ 4.6 Analysis of Contextual and Parametric Knowledge Conflict and Fusion
When contextual knowledge and parametric knowledge conflict, LLMs may struggle to determine which source of information is more reliable [39; 48]. A key distinction between parametric RAG and standard RAG lies in whether external documents are used as contextual knowledge or transforming into parametric knowledge. Additionally, we aim to explore why combining parameter injection with in-context injection results in superior performance.
@@
Vanilla represents the answer from original LLMs without any external knowledge. As shown in Table 4, the vanilla LLM contains incorrect parametric knowledge. While standard RAG and DyPRAG address this by performing in-context injection and parameter injection, respectively. Both approaches introduce hallucinations that result in incorrect answers. In contrast, our DyPRAG-Combine effectively integrates contextual knowledge with transformed parametric knowledge, enabling it to provide correct answers and demonstrating its ability to leverage both types of knowledge effectively.
Compared to standard RAG, DyPRAG-Combine transforms the retrieved documents into parametric knowledge before injecting them into the model. This approach ensures that the LLM already contains relevant knowledge when answering questions, unlike standard RAG which often suffers from well-known hallucination issues (i.e., knowledge conflicts) [39; 37]. As a result, DyPRAG-Combine offers a practical and efficient solution to directly relieve the common hallucination problem in standard RAG, providing a plug-and-play improvement to enhance RAG’s performance in real-world applications.
@@
The retrieved documents in our experiments are primarily sourced from Wikipedia, which are already encountered by LLMs during their pre-training phase. In this section, we further investigate how DyPRAG performs on unseen contextual knowledge using the RAGTruth benchmark [45]. Specifically, we randomly sample 100 examples from the QA-type sub-dataset, which presents greater challenges (e.g., the required answers are only accessible in carefully crafted context). As shown in Figure 4, our DyPRAG-Combine significantly outperforms standard RAG, as evaluated by GPT-4o. This demonstrates that DyPRAG effectively enables LLMs to better internalize contextual knowledge and mitigate hallucinations, even when handling unseen data.  [^1] We provide further detail analysis of contextual and parametric knowledge in Appendix D.
4.7 Ablation Study
Effect of Intermediate Dimension $p$.
Figure 5: Ablation study of intermediate dimension $p$ of $\mathcal{F}^{\prime}_{\phi}$. The backbone model is the Qwen2.5-1.5B. The inference time is computed by average time of CWQ with batch_size of 1. The encode time is highlighted in red, while the translate time is marked in blue.
Method
CWQ
F1
Inference Time
Storage Cost
RAG 28.32 1.20 (1x) -
PRAG 30.82 0.56 (0.47x) 672 (1x)
DyPRAG ($p=2$) 32.66 0.56+0.13+0.060 (0.625x) 7.71 (0.011x)
DyPRAG ($p=4$) 33.26 0.56+0.13+0.062 (0.627x) 15.42 (0.023x)
DyPRAG ($p=16$) 32.08 0.56+0.13+0.055 (0.621x) 61.70 (0.092x)
DyPRAG ($p=32$) 31.94 0.56+0.13+0.060 (0.625x) 123.39 (0.184x)
Figure 6: Ablation study of alignment loss. The backbone model is the LLaMA3.2-1B.
Methods HQA PQA CWQ
Bridge Compare Total
DyPRAG 14.05 43.9 19.97 8.37 36.86
Ablation Study
w/o $\mathcal{L}_{kl}$ 7.35 35.28 13.12 1.93 22.85
w/o $\mathcal{L}_{mse}$ 12.78 42.11 17.65 5.94 32.98
w/o $\mathcal{L}_{kl},\mathcal{L}_{mse}$ 12.66 39.46 18.26 2.42 34.92
Effect of Alignment Loss.
@@ Appendix A Detailed Cost Comparison
Additionally, to train our $\mathcal{F}^{\prime}_{\phi}$ for $E_{2}$ epochs, we need to perform both forward and backward passes (the backward pass requires twice the cost of the forward pass) on one QA pair and its corresponding document in each step. This results in a cost of $N\times E_{2}\times 9(|qa|+|d|)^{2}\times\text{ATTN}+3(|qa|+|d|)\times\text{FFN}$, with a negligible cost for translation. As shown in Figure 8 and  8, our DyPRAG achieves stable results with as few as 480 examples (even fewer is powerful), while $M=3000$ in our experiments, and this value would be significantly larger in real-world applications.
@@ B.1 Implementation Details
All experiments use the publicly available Hugging Face implementations of LLaMA and Qwen. To ensure fairness, DyPRAG and all baselines share the same prompt template following [36](in Figure  15 and  16) and adopt of greedy decoding for result reproducibility.
@@
Motivated by [23], we use simple MLP hypernetwork to transform embedding into adapter parameters. Through cross validation, the learning rate was set to $1\times 10^{-5}$, and the training epoch was set to 1 which making the overall alignment process quickly. The truncation max length of text is set to 3000, which is larger than most retrieved documents.
@@
For each dataset, we select the first 300 examples for testing and evaluate performance using F1 score for IIRC and Accuracy for SQA as metrics. For IIRC, we adopt the few-shot prompts from [35], while SQA is evaluated in a zero-shot setting. Notably, the same prompt format (in Figure 15 and  16) from the main experiment is used to ensure a fair comparison.
@@
DyPRAG (ours)  25.18  19.67 22.43
DyPRAG-Combine (ours)  26.46  27.57 27.02
@@ Appendix C Supplement Experiment Results
We adjust the pre-selected size of the training dataset composed of Doc-Param pairs, increasing it from 480 to 4800. As shown in Figure 8 and 8, DyPRAG achieves strong performance even with just 480 training examples. The performance remains remarkably stable across different dataset sizes, indicating that our design, $\mathcal{F}^{\prime}_{\phi}$, is capable of learning the underlying mapping between documents and parameters with minimal data.
Table 6: Supplement ablation study of alignment loss. The backbone model is the LLaMA3.2-1B.
Methods 2WQA
Compare Bridge Inference Compose Total
DyPRAG 0.5125 0.4815 0.1735 0.0754 0.2531
@@
w/o $\mathcal{L}_{kl}$ 0.2954 0.3774 0.1294 0.0578 0.2027
w/o $\mathcal{L}_{mse}$ 0.5606 0.3696 0.1729 0.084 0.2728
w/o $\mathcal{L}_{kl},\mathcal{L}_{mse}$ 0.4523 0.2484 0.1674 0.0748 0.2343
Effect of Alignment Loss in 2WQA.
In Section 4.7, we have already investigated the effect of combining different components of the alignment loss. As shown in Table 6, the trend is consistent with Table 6. Our DyPRAG achieves the highest average score of 29.92%, while the model without $\mathcal{L}_{\text{mse}}$ achieves the second-best score of 29.20%. These experimental results demonstrate that each element of the alignment loss contributes to effectively training $\mathcal{F}^{\prime}_{\phi}$ to generate $p^{\prime}$.
@@
Appendix D Further Analysis of Contextual and Parametric Knowledge Conflict
@@
Appendix E Visualization of Parameter Translator Workflow.
@@
Appendix F Prompt for Main Experiments Evaluation
In the main experiments, we used the following prompt to assess the performance of DyPRAG and other baseline models in Figure 15 and  16:
@@ Appendix F Prompt for Main Experiments Evaluation
Appendix G Prompt for Knowledge Internalization Evaluation

### v2

@@ 4.6 Analysis of Contextual and Parametric Knowledge Conflict and Fusion
As shown in Table 4, the vanilla LLM contains incorrect parametric knowledge. While standard RAG and DyPRAG try to address this by performing in-context injection and parameter injection, respectively. Both approaches introduce hallucinations that result in incorrect answers.
In contrast, DyPRAG-Combine effectively integrates contextual knowledge with transformed parametric knowledge, enabling it to provide correct answers and demonstrating its ability to leverage both types of knowledge effectively. Compared to RAG, DyPRAG-Combine transforms the retrieved documents into parametric knowledge before injecting them into the model. This approach ensures that the LLM already contains relevant knowledge when answering questions, unlike RAG which often suffers from well-known hallucination issues [42; 40]. As a result, DyPRAG-Combine offers a practical and efficient solution to directly relieve the common hallucination problem in standard RAG, providing a plug-and-play improvement to enhance RAG’s performance in real-world applications.
@@
The retrieved documents in our experiments are primarily sourced from Wikipedia, which are already encountered by LLMs during the pre-training phase. In this section, we further investigate how DyPRAG performs on unseen contextual knowledge using the RAGTruth benchmark [33]. Specifically, we randomly sample 100 examples from the QA-type sub-dataset, which presents greater challenges (e.g., the required answers are only accessible in carefully crafted context). As shown in Figure 3, DyPRAG-Combine significantly outperforms RAG. This demonstrates that DyPRAG effectively enables LLMs to better internalize contextual knowledge and mitigate hallucinations, even when handling unseen data.  [^1] We provide further detail analysis of contextual and parametric knowledge in Appendix E.
@@ Appendix A Detailed Cost Comparison
Additionally, to train our $\mathcal{F}^{\prime}_{\phi}$ for $E_{2}$ epochs, we need to perform both forward and backward passes (the backward pass requires twice the cost of the forward pass) on one QA pair and its corresponding document in each step. This results in a cost of $N\times E_{2}\times 9(|qa|+|d|)^{2}\times\text{ATTN}+3(|qa|+|d|)\times\text{FFN}$, with a negligible cost for translation. As shown in Figure 6 and  6, our DyPRAG achieves stable results with as few as 480 examples (even fewer is powerful), while $M=3000$ in our experiments, and this value would be significantly larger in real-world applications.
@@ B.1 Implementation Details
QA Datasets.
To ensure a comprehensive evaluation, we assess our method using the following datasets:
@@
All experiments use the publicly available Hugging Face implementations of LLaMA and Qwen. To ensure fairness, DyPRAG and all baselines share the same prompt template in Figure  14 and  15 following [39] and adopt of greedy decoding for result reproducibility. The max number of new tokens is set to 128.
@@
Motivated by [25], we use simple MLP hypernetwork to transform embedding into adapter parameters. Through cross validation, the learning rate was set to $1\times 10^{-5}$, and the training epoch was set to 1 which making the overall alignment process quickly. The truncation max length of text is set to 3000, which is larger than most retrieved documents. The performance reports for Qwen2.5-1.5B and LLaMA3.2-1B in Table 1 are based on training with 4,800 examples, while LLaMA3-8B is trained on 2,400 examples (except for 480 examples on 2WQA).
@@
For each dataset, we select the first 300 examples for testing and evaluate performance using F1 score for IIRC and Accuracy for SQA as metrics. Both datasets provide with ground-truth passages which indicate a more rigorous evaluation setting. For IIRC, we adopt the few-shot prompts from [38], while SQA is evaluated in a zero-shot setting. Notably, the same prompt format (in Figure 14 and  15) from the main experiment is used to ensure a fair comparison.
@@
DyPRAG (ours)  26.46  19.67 23.07
DyPRAG-Combine (ours)  25.18  27.57 26.38
@@ Appendix C Supplement Experiment Results
Varying Number of Injected Documents
For standard RAG, the number of retrieved documents, denoted as $c$, is a crucial hyperparameter to tune. Recent studies [23, 47] have investigated the impact of longer context lengths on standard RAG. However, the effect of the number of injected documents in parametric form remains underexplored. Our proposed DyPRAG framework can seamlessly adapt to this scenario due to its inherent flexibility.
@@
We adjust the pre-selected size of the training dataset composed of Doc-Param pairs, increasing it from 480 to 4800. As shown in Figure 6 and 6, DyPRAG achieves strong performance even with just 480 training examples. The performance remains remarkably stable across different dataset sizes, indicating that our design, $\mathcal{F}^{\prime}_{\phi}$, is capable of learning the underlying mapping between documents and parameters with minimal data.
Table 7: Ablation study of alignment loss. The backbone model is the LLaMA3.2-1B.
Methods 2WQA HQA PQA CWQ Avg
Compare Bridge Inference Compose Total Bridge Compare Total
DyPRAG 51.25 48.15 17.35 7.54 25.31 14.05 43.9 19.97 8.37 36.86 25.28
@@
w/o $\mathcal{L}_{kl}$ 29.54 37.74 12.94 5.78 20.27 7.35 35.28 13.12 1.93 22.85 18.68
w/o $\mathcal{L}_{mse}$ 56.06 36.96 17.29 8.40 27.28 12.78 42.11 17.65 5.94 32.98 23.38
w/o $\mathcal{L}_{kl},\mathcal{L}_{mse}$ 45.23 24.84 16.74 7.48 23.43 12.66 39.46 18.26 2.42 34.92 22.54
Effect of Alignment Loss.
@@
Can Perplexity Reflects Knowledge Conflict?
Recent studies have explored methods to detect hallucinations in LLMs and RAG systems by leveraging various metrics [6, 40]. Among these, we adopt the simplest yet effective metric, Perplexity (PPL) [35], to evaluate knowledge conflicts. As illustrated in Figure8, Vanilla and DyPRAG exhibit higher PPL, while DyPRAG-Combine and RAG demonstrate significantly lower PPL. However, these results are inconsistent with the findings in Table 1 and Table 3. For instance, although DyPRAG-Combine achieves the best performance on IIRC, its calculated PPL suggests a higher probability of knowledge conflicts, which is clearly incorrect.
We hypothesize that this discrepancy arises primarily from variations in model parameters caused by parameter injection in DyPRAG, which is undetectable with the current evaluation method. Developing more advanced methods to detect knowledge conflicts in DyPRAG remains an interesting avenue for future research.
@@
Figure 8: Comparison of average PPL. Smaller PPL means less conflict. The backbone model is the LLaMA3.2-1B.
Appendix D Why Vanilla Outperforms RAG Occasionally ?
In this section, we provide a detailed analysis of why the vanilla model occasionally outperforms RAG. As shown in Table 1, the vanilla model surpasses RAG most significantly in 2WQA, although the results vary across different models. For instance, the vanilla model outperforms RAG by 2.62 and 0.99 in Qwen2.5-1.5B and LLaMA3-8B on average, respectively. After analyzing the cases, we identify two key issues that affect RAG’s performance: 1) Poor Retriever. Following [39], we use BM25 as the retriever. However, in many cases, the retrieved documents contain only similar words rather than relevant content. This results in the provided content being unhelpful or even detrimental to the performance of LLMs. 2) Already Seen Data. During the pre-training stages of the selected LLMs [49, 31, 30], the external source we use (i.e., Wikipedia) has already been seen. This allows LLMs to answer certain questions independently, especially in simpler tasks like 2WQA. Moreover, the inclusion of incorrect or irrelevant context further degrades the performance of LLMs, as observed in Table 1.
A more rigorous evaluation setting should include ground-truth passages and ensure no or less data leakage.Under this setting, as shown in Table 3, the performance of the vanilla model is significantly lower than that of RAG, which aligns with our hypothesis. For instance, the vanilla model achieves only 8.78% and 1.00% accuracy on Qwen2.5-1.5B for IIRC and SQA, respectively. In contrast, DyPRAG demonstrates a notable improvement in test-time knowledge, achieving 10.23% and 15.67% accuracy on Qwen2.5-1.5B for IIRC and SQA, respectively. These results underscore the critical role of RAG while showcasing the ability of our proposed DyPRAG to seamlessly enhance OOD knowledge effectively. Furthermore, DyPRAG-Combine establishes a superior RAG paradigm by delivering even better performance under these more challenging conditions. In summary, we believe that this more rigorous experimental setting better validates our proposed method.
Appendix E Further Analysis of Contextual and Parametric Knowledge Conflict
@@
Appendix F Visualization of Parameter Translator Workflow.
@@
Appendix G Prompt for Main Experiments Evaluation
In the main experiments, we used the following prompt to assess the performance of DyPRAG and other baseline models in Figure 14 and  15:
@@ Appendix G Prompt for Main Experiments Evaluation
Appendix H Prompt for Knowledge Internalization Evaluation